# Continual Learning for Image Captioning through Improved Image-Text Alignment

### v1

@@ 1 Introduction
Understanding how to generate natural language descriptions of images is a fundamental problem at the intersection of computer vision and natural language processing. Recent advances in deep learning have enabled remarkable progress in static image captioning [1], where large-scale vision-language models are trained to describe images with fluent, contextually appropriate sentences. However, these conventional systems are designed for fixed, closed-world scenarios in which all training data are available upfront, and the models are not required to adapt to new visual categories or linguistic concepts over time. In contrast, continual learning is an emerging research paradigm that seeks to endow captioning systems with the ability to learn from a stream of data composed of sequentially arriving tasks, new domains, or novel object categories, all without revisiting previous data [2, 3]. This setting introduces unique challenges associated with lifelong learning, such as the risk of catastrophic forgetting [4], in which previously acquired knowledge is overwritten as the model adapts to novel information. This paradigm can be applied to the case of image captioning, [5, 6]. Successful continual image captioning systems would be capable of incrementally expanding their understanding and generating accurate, coherent captions for both seen and unseen concepts as they are exposed to them. The motivation for continual image captioning extends beyond academic interest. It is highly relevant for real-world applications such as autonomous agents, assistive devices, and interactive robotics, which operate in dynamic environments and must continually assimilate new visual and linguistic patterns [7]. Unlike classification, captioning is a generative task that demands cross-modal reasoning—grounding visual information in language with both fine-grained perceptual detail and pragmatic contextualization. As such, continual image captioning provides a rich testbed for studying not only incremental learning and memory-efficient architectures, but also the complex interplay between vision and language adaptation [8]. It invites exploration of new training strategies, self-supervised learning, and regularization schemes specifically tailored for vision-language models in an open world. Despite growing interest, continual learning methods for computer vision have been largely explored in the context of image classification [3], with comparatively less attention to more complex vision-language generation settings. Prior attempts to mitigate catastrophic forgetting using prompt-based methods or contrastive learning have shown preliminary promise in related tasks [9, 10], but unique challenges arise for generative modeling: the need for preserving object-word alignments, maintaining linguistic diversity, and avoiding loss of semantic granularity across tasks. In this work, we address these challenges by introducing a novel multi-loss training framework for continual image captioning. Building on a pretrained ViT-GPT-2 architecture [11], our approach combines the standard cross-entropy loss with (1) a prompt-based cosine similarity loss ($L_{\text{nouns}}$) to align image embeddings with semantic prompt representations, (2) a CLIP-based cosine similarity loss ($L_{\text{CLIP}}$) to align image embeddings to target caption textual embedding. The framework is initiated with prompt-based, later switching to caption-based alignment training. (3) A language-guided contrastive loss on the class level ($L_{\mathrm{LGCL}}$), similar to [12], which promotes semantic discriminability across tasks through a triplet-based cosine similarity scheme. Our proposed dynamic loss balancing mechanism further prevents any single loss component from dominating training, reducing the risk of overfitting or under-adaptation. All four losses are summed together into a single loss, which is used for backpropagation. Through experiments on a continual split of the MS-COCO dataset [6], we show that our method outperforms baseline and state-of-the-art approaches, in particular exhibiting improved semantic retention. Note that our method incurs no inference-time overhead, making it attractive for practitioners in resource-constrained environments. Our contributions can be summarized as follows.
@@ 2 Related Work
Image captioning, the task of generating descriptive natural language sentences for images, has seen remarkable advances through deep learning. Early approaches, such as the Neural Image Caption Generator (NIC) [13], demonstrated promising results and inspired a series of works aimed at improving accuracy and expressiveness [14]. Despite progress, these models are typically trained offline on fixed datasets and crucially suffer from catastrophic forgetting when new tasks are encountered sequentially, a phenomenon well documented in the continual learning literature [4], [3]. To address catastrophic forgetting, continual learning methods have emerged and can be largely classified into the following categories of approaches: regularization-based, replay-based, optimization-based, representation-based, architecture-based [3]. Prompt-based continual learning is a recent approach that belongs to the category of representation-based approaches and introduces parameter-efficient solutions by leveraging pre-trained models and learnable prompts, synthesizing transfer learning ideas from natural language processing (NLP) and extending them to vision-language models [9], [10], [12]. This approach bypasses the need to store or replay task data and can encode task information directly in prompts without requiring explicit task IDs [15]. Recent work integrating language guidance into vision models has demonstrated improved generalization, especially for unseen classes, by aligning visual and semantic spaces [16], [17], [18]. Within image captioning, the challenge of continual learning is rarely studied. Methods such as ContCap [6] employed freezing, pseudo-labeling, and feature distillation. The approaches [5], [19] focussed particularly on recurrent approaches. Yet prompt-based approaches in image captioning remain largely unexplored. Our research advances this frontier by integrating prompt-based continual learning and language guidance, establishing a novel, efficient paradigm for incremental image captioning that addresses catastrophic forgetting without rehearsal or extensive retraining.
@@ 3.1 Notation
Continual Learning focuses on training a machine learning model on a data stream originating from a sequence of tasks. Let us denote this sequence of tasks as $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{T}\}$, where each task $D_{t}$ consists of tuples $(x_{tk},y_{tk})$. Here, $x_{tk}\in\mathcal{X}$ represents the input image $k$ in task $t$ in the RGB color space, and $y_{tk}\in\mathcal{Y}$ denotes the corresponding label associated with task $t$. For the vocabulary we utilize the pretrained GPT-2 tokenizer [20], with a vocabulary of $50,257$ tokens. This vocabulary, originally trained on a WebText corpus, allows for efficient and consistent tokenization of textual inputs without modifying the underlying language model. By leveraging this fixed vocabulary, our method ensures compatibility with the GPT-2 architecture.
@@ 3.2 Proposed Approach
Language-Guided Contrastive Loss ($L_{\mathrm{LGCL}}$) Following recent advances in language guidance for prompt-based continual learning [12], we further enhance the model’s discriminative capacity by employing a language-guided contrastive objective. Here, we draw inspiration from triplet loss formulations [21], aiming to maximize the alignment between semantically corresponding image-text pairs while repelling mismatching associations.
@@
This objective encourages $\mathbf{v(\theta)}_{\text{img}}$ to remain proximate to its correct linguistic counterpart, while maintaining a margin from negatives.
@@
4 Experiments
4.1 Metrics
We evaluated the performance of our models using standard image captioning metrics, including BLEU [22], ROUGE [23], CIDEr [24], and METEOR [25], used in COCO-Caption. Note, METEOR tracks semantic consistency particularly well at the sentence level, because it aligns hypotheses to references using stems, synonyms as well as paraphrase tables and balances precision and recall [26]. We also report CLIPScore [8], which evaluates caption quality by measuring the cosine similarity between the CLIP image and text embeddings and correlates very well with human judgment[8]. In all cases, higher scores indicate better performance. To quantify forgetting, we consider an end-of-task forgetting metric and compute the relative change in metric performance for each task as follows:
[…]
This expresses the proportion of performance lost due to forgetting.
4.2 Experimental Setup
All experiments were performed on a system equipped with an NVIDIA GeForce RTX 4060 Ti GPU. For both ContCap and RATT dataset splits, the setup remained the same. We trained the model for $5$ epochs, with the batch size of 32 and a learning rate (LR) of 1e-5. We used the AdamW optimizer[27], with default momentum parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a base learning rate of $\eta_{0}=\text{LR}/25$, where LR is the peak learning rate.
4.3 Comparison to State-of-the-Art Methods
Table 1, in the same tabular format as in the original paper [6] (Table III), shows that our CLICITA model achieves substantial improvements over the best ContCap baselines ($S_{19}$ and $S_{multiple}$) in nearly all key evaluation metrics. The most notable gains are observed in CIDEr, METEOR, and BLEU-4, indicating that CLICITA significantly improves both the relevance and the fluency of generated captions. While ROUGE-L lags slightly behind the best ContCap variant.
Table 1: Comparison of Best ContCap ($S_{19}$ and $S_{multiple}$) vs. CLICITA
BLEU-1 47.1 ($FD$) 53.6 ($FD$) 67.10 +20.00 +13.50
BLEU-4 6.6 ($P$) 10.5 ($P$) 22.00 +15.40 +11.50
ROUGE-L 34.0 ($FD$) 40.0 ($FD$) 36.10 +2.10 -3.90
METEOR 11.2 ($D_{F}$) 14.5 ($E_{F}$) 27.90 +16.70 +13.40
CIDEr 10.0 ($P$) 19.2 ($E_{F}$) 66.20 +56.20 +47.00
Table 2 presents a task-wise comparison on the RATT split, in similar tabular format as in the [5], (Table 2). While CLICITA reports lower BLEU-4 and CIDEr scores across all tasks, indicating a slight drop in n-gram precision and content specificity, it consistently outperforms RATT on METEOR, a key semantic metric. This suggests that CLICITA produces captions with better overall meaning alignment and linguistic fluency, even when exact n-gram matches are fewer.
Metric Transport Animals Sports Food Interior
In the following sections we evaluated our method on two different datasets in two distinct settings to investigate the improvement of the designed mechanism over the baseline pretrained model. Note that pre-trained models already serve the continual learning setting due to flat minima more often found in pre-trained models [3]. The models are:
4.4 Experimental Result - ContCap Dataset Split
Following the training approach outlined in ContCap [6], we incrementally introduced five object classes. These five classes (tasks) were sequentially added one by one into our model: person, sports ball, tv, toilet, and bottle, with the class ’bottle’ being the final task trained for the model. The focus is on knowledge retention. Table 3 shows the average forgetting scores for this dataset.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (ContCap) -0.32 -2.42 -0.96 2.66 -4.42 -0.44 -5.9
4.5 Experimental Results - RATT dataset split
In RATT [5] split they define five distinct tasks (Transport, Animals, Sports, Food, Interior) based on object categories and processes the dataset to ensure non-overlapping image assignments across tasks. The implementation strictly enforces the paper’s reported image counts per task (14,266/3,431/3,431 for Transport, 9,314/2,273/2,273 for Animals, etc.) by first filtering images containing relevant categories, removing duplicates across tasks, and maintaining only images with more or equal to 5 captions and keeping only the first 5 captions. The validation set is further split 50/50 into validation and test sets. We compare our CLICITA method against Pre-Trained Basemodel, again with a focus is on knowledge retention.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (RATT) -0.14 -2.50 -1.36 0.24 -3.62 -0.68 -7.06
5 Discussion
6 Conclusion
Acknowledgments

### v2

@@ 1 Introduction
Understanding how to generate natural language descriptions of images is a fundamental problem at the intersection of computer vision and natural language processing. Recent advances in deep learning have enabled remarkable progress in static image captioning [xu2015show], where large-scale vision-language models are trained to describe images with fluent, contextually appropriate sentences. However, these conventional systems are designed for fixed, closed-world scenarios in which all training data are available upfront, and the models are not required to adapt to new visual categories or linguistic concepts over time. In contrast, continual learning is an emerging research paradigm that seeks to endow captioning systems with the ability to learn from a stream of data composed of sequentially arriving tasks, new domains, or novel object categories, all without revisiting previous data [parisi2019continual, wang_comprehensive_2024]. This setting introduces unique challenges associated with lifelong learning, such as the risk of catastrophic forgetting [mccloskey_catastrophic_1989], in which previously acquired knowledge is overwritten as the model adapts to novel information. This paradigm can be applied to the case of image captioning, [chiaro2020ratt, nguyen_contcap_2020]. Successful continual image captioning systems would be capable of incrementally expanding their understanding and generating accurate, coherent captions for both seen and unseen concepts as they are exposed to them. The motivation for continual image captioning extends beyond academic interest. It is highly relevant for real-world applications such as autonomous agents, assistive devices, and interactive robotics, which operate in dynamic environments and must continually assimilate new visual and linguistic patterns [golroudbari_recent_2023]. Unlike classification, captioning is a generative task that demands cross-modal reasoning—grounding visual information in language with both fine-grained perceptual detail and pragmatic contextualization. As such, continual image captioning provides a rich testbed for studying not only incremental learning and memory-efficient architectures, but also the complex interplay between vision and language adaptation [hessel2021clipscore]. It invites exploration of new training strategies, self-supervised learning, and regularization schemes specifically tailored for vision-language models in an open world. Despite growing interest, continual learning methods for computer vision have been largely explored in the context of image classification [wang_comprehensive_2024], with comparatively less attention to more complex vision-language generation settings. Prior attempts to mitigate catastrophic forgetting using prompt-based methods or contrastive learning have shown preliminary promise in related tasks [Wang2022, wang2022dualprompt], but unique challenges arise for generative modeling: the need for preserving object-word alignments, maintaining linguistic diversity, and avoiding loss of semantic granularity across tasks. In this work, we address these challenges by introducing a novel multi-loss training framework for continual image captioning. Building on a pretrained ViT-GPT-2 architecture [shreydan2023visiongpt2], our approach combines the standard cross-entropy loss with (1) a prompt-based cosine similarity loss ($L_{\text{nouns}}$) to align image embeddings with semantic prompt representations, (2) a CLIP-based cosine similarity loss ($L_{\text{CLIP}}$) to align image embeddings to target caption textual embedding. The framework is initiated with prompt-based, later switching to caption-based alignment training. (3) A language-guided contrastive loss on the class level ($L_{\mathrm{LGCL}}$), similar to [khan2023introducinglanguageguidancepromptbased], which promotes semantic discriminability across tasks through a triplet-based cosine similarity scheme. Our proposed dynamic loss balancing mechanism further prevents any single loss component from dominating training, reducing the risk of overfitting or under-adaptation. All four losses are summed together into a single loss, which is used for backpropagation. Through experiments on a continual split of the MS-COCO dataset [nguyen_contcap_2020], we show that our method outperforms baseline and state-of-the-art approaches, in particular exhibiting improved semantic retention. Note that our method incurs no inference-time overhead, making it attractive for practitioners in resource-constrained environments. Our contributions can be summarized as follows.
@@ 2 Related Work
Image captioning, the task of generating descriptive natural language sentences for images, has seen remarkable advances through deep learning. Early approaches, such as the Neural Image Caption Generator (NIC) [vinyals_show_2015], demonstrated promising results and inspired a series of works aimed at improving accuracy and expressiveness [cornia_show_2019]. Despite progress, these models are typically trained offline on fixed datasets and crucially suffer from catastrophic forgetting when new tasks are encountered sequentially, a phenomenon well documented in the continual learning literature [mccloskey_catastrophic_1989], [wang_comprehensive_2024]. To address catastrophic forgetting, continual learning methods have emerged and can be largely classified into the following categories of approaches: regularization-based, replay-based, optimization-based, representation-based, architecture-based [wang_comprehensive_2024]. Prompt-based continual learning is a recent approach that belongs to the category of representation-based approaches and introduces parameter-efficient solutions by leveraging pre-trained models and learnable prompts, synthesizing transfer learning ideas from natural language processing (NLP) and extending them to vision-language models [Wang2022], [wang2022dualprompt], [khan2023introducinglanguageguidancepromptbased]. This approach bypasses the need to store or replay task data and can encode task information directly in prompts without requiring explicit task IDs [pfeiffer_adapterfusion_2021]. Recent work integrating language guidance into vision models has demonstrated improved generalization, especially for unseen classes, by aligning visual and semantic spaces [radford_learning_2021], [naeem_i2mvformer_2022], [naeem_learning_2021]. Within image captioning, the challenge of continual learning is rarely studied. Methods such as ContCap [nguyen_contcap_2020] employed freezing, pseudo-labeling, and feature distillation. The approaches [chiaro2020ratt], [li_incremental_2024] focussed particularly on recurrent approaches. Yet prompt-based approaches in image captioning remain largely unexplored. Our research advances this frontier by integrating prompt-based continual learning and language guidance, establishing a novel, efficient paradigm for incremental image captioning that addresses catastrophic forgetting without rehearsal or extensive retraining.
@@ 3.1 Notation
Continual Learning focuses on training a machine learning model on a data stream originating from a sequence of tasks. Let us denote this sequence of tasks as $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{T}\}$, where each task $D_{t}$ consists of tuples $(x_{tk},y_{tk})$. Here, $x_{tk}\in\mathcal{X}$ represents the input image $k$ in task $t$ in the RGB color space, and $y_{tk}\in\mathcal{Y}$ denotes the corresponding label associated with task $t$. For the vocabulary we utilize the pretrained GPT-2 tokenizer [radford2019language], with a vocabulary of $50,257$ tokens. This vocabulary, originally trained on a WebText corpus, allows for efficient and consistent tokenization of textual inputs without modifying the underlying language model. By leveraging this fixed vocabulary, our method ensures compatibility with the GPT-2 architecture.
@@ 3.2 Proposed Approach
Language-Guided Contrastive Loss ($L_{\mathrm{LGCL}}$) Following recent advances in language guidance for prompt-based continual learning [khan2023introducinglanguageguidancepromptbased], we further enhance the model’s discriminative capacity by employing a language-guided contrastive objective. Here, we draw inspiration from triplet loss formulations [feragen2015], aiming to maximize the alignment between semantically corresponding image-text pairs while repelling mismatching associations.
@@
This objective encourages $\mathbf{v(\theta)}_{\text{img}}$ to remain proximate to its correct linguistic counterpart, while maintaining a margin from negatives. The pseudocode for the algorithm can be found in the appendix A.
4 Experiments
4.1 Metrics
We evaluated the performance of our models using standard image captioning metrics, including BLEU [papineni_bleu_2001], ROUGE [lin_rouge_2004], CIDEr [vedantam_cider_2015], and METEOR [lavie_meteor_2007], used in COCO-Caption. Note, METEOR tracks semantic consistency particularly well at the sentence level, because it aligns hypotheses to references using stems, synonyms as well as paraphrase tables and balances precision and recall [DBLP:journals/corr/ChenFLVGDZ15]. We also report CLIPScore [hessel2021clipscore], which evaluates caption quality by measuring the cosine similarity between the CLIP image and text embeddings and correlates very well with human judgment[hessel2021clipscore]. In all cases, higher scores indicate better performance. To quantify forgetting, we consider an end-of-task forgetting metric and compute the relative change in metric performance for each task as follows:
[…]
This expresses the proportion of performance lost due to forgetting.
4.2 Experimental Setup
All experiments were performed on a system equipped with an NVIDIA GeForce RTX 4060 Ti GPU. For both ContCap and RATT dataset splits, the setup remained the same. We trained the model for $5$ epochs, with the batch size of 32 and a learning rate (LR) of 1e-5. We used the AdamW optimizer[loshchilov2019decoupledweightdecayregularization], with default momentum parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a base learning rate of $\eta_{0}=\text{LR}/25$, where LR is the peak learning rate.
4.3 Comparison to State-of-the-Art Methods
Table 1, in the same tabular format as in the original paper [nguyen_contcap_2020] (Table III), shows that our CLICITA model achieves substantial improvements over the best ContCap baselines ($S_{19}$ and $S_{multiple}$) in nearly all key evaluation metrics. The most notable gains are observed in CIDEr, METEOR, and BLEU-4, indicating that CLICITA significantly improves both the relevance and the fluency of generated captions. While ROUGE-L lags slightly behind the best ContCap variant.
Table 1: Comparison of Best ContCap ($S_{19}$ and $S_{multiple}$) vs. CLICITA
BLEU-1 47.1 ($FD$) 53.6 ($FD$) 67.10 +20.00 +13.50
BLEU-4 6.6 ($P$) 10.5 ($P$) 22.00 +15.40 +11.50
ROUGE-L 34.0 ($FD$) 40.0 ($FD$) 36.10 +2.10 -3.90
METEOR 11.2 ($D_{F}$) 14.5 ($E_{F}$) 27.90 +16.70 +13.40
CIDEr 10.0 ($P$) 19.2 ($E_{F}$) 66.20 +56.20 +47.00
Table 2 presents a task-wise comparison on the RATT split, in similar tabular format as in the [chiaro2020ratt], (Table 2). While CLICITA reports lower BLEU-4 and CIDEr scores across all tasks, indicating a slight drop in n-gram precision and content specificity, it consistently outperforms RATT on METEOR, a key semantic metric. This suggests that CLICITA produces captions with better overall meaning alignment and linguistic fluency, even when exact n-gram matches are fewer.
Metric Transport Animals Sports Food Interior
In the following sections we evaluated our method on two different datasets in two distinct settings to investigate the improvement of the designed mechanism over the baseline pretrained model. Note that pre-trained models already serve the continual learning setting due to flat minima more often found in pre-trained models [wang_comprehensive_2024]. The models are:
4.4 Experimental Result - ContCap Dataset Split
Following the training approach outlined in ContCap [nguyen_contcap_2020], we incrementally introduced five object classes. These five classes (tasks) were sequentially added one by one into our model: person, sports ball, tv, toilet, and bottle, with the class ’bottle’ being the final task trained for the model. The focus is on knowledge retention. Table 3 shows the average forgetting scores for this dataset.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (ContCap) -0.32 -2.42 -0.96 2.66 -4.42 -0.44 -5.9
4.5 Experimental Results - RATT dataset split
In RATT [chiaro2020ratt] split they define five distinct tasks (Transport, Animals, Sports, Food, Interior) based on object categories and processes the dataset to ensure non-overlapping image assignments across tasks. The implementation strictly enforces the paper’s reported image counts per task (14,266/3,431/3,431 for Transport, 9,314/2,273/2,273 for Animals, etc.) by first filtering images containing relevant categories, removing duplicates across tasks, and maintaining only images with more or equal to 5 captions and keeping only the first 5 captions. The validation set is further split 50/50 into validation and test sets. We compare our CLICITA method against Pre-Trained Basemodel, again with a focus is on knowledge retention.
Split BLEU-1 BLEU-4 ROUGE-L METEOR CIDEr CLIP Average Forgetting
Average Forgetting (%)
CLICITA (RATT) -0.14 -2.50 -1.36 0.24 -3.62 -0.68 -7.06
5 Discussion
6 Conclusion
Acknowledgments
Appendix A Algorithms